OpenStudy (anonymous):

If this is true... [If a function is differentiable at x=a then f in continous at x=a.] Then why is this false? [If a function f is continous at x=a then f is differentiable at x=a.]

jimthompson5910 (jim_thompson5910):

If a function f is continous at x=a then f is differentiable at x=a is false look at the absolute value function

jimthompson5910 (jim_thompson5910):

it's continuous at x = 0 for f(x) = |x| but it's not differentiable there

OpenStudy (anonymous):

Then how is the first one true then?.. it's just opposite?

jimthompson5910 (jim_thompson5910):

If a function is differentiable at x=a then f in continuous at x=a is true because the whole notion of differentiability is based on limits and continuity

OpenStudy (anonymous):

So, if its differentiable then it's continuous. But, if its continuous then it's not differentiable.

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

first one is always guaranteed to be true, the second isn't always true (so it's false in general)

OpenStudy (anonymous):

haha what in the world....

jimthompson5910 (jim_thompson5910):

and you can show the second isn't always true using the absolute value function as an example

jimthompson5910 (jim_thompson5910):

hmm what do you mean?

OpenStudy (anonymous):

They are not logically equivalent statements. If A is true, then B is true does not imply if B is true, then A is true. A simple example: If a car goes 70km/h then it receives a speeding ticket. If a car receives a speeding ticket, it is going 70km/h. If a function is differentiable, it is continuous. If a function is continuous it may or may not be differentiable.

OpenStudy (anonymous):

^ that makes a lot of sense!

OpenStudy (anonymous):

I was think that if the first statement was true, then the second had to be true.. But, I guess not.

OpenStudy (anonymous):

:) yeah it took me awhile to sort out the formal logic they throw at you in Calculus...contrapositives and all that

OpenStudy (anonymous):

the reason that if a function is differentiable at \(a\) then it must be continuous at \(a\) comes from the the following: if \(f\) is differentiable at \(a\) it means \[\lim_{h\to 0}\frac{f(a+h)-f(a)}{h}\] exists, i.e the limit exists

OpenStudy (anonymous):

now in the above limit the denominator is going to zero this means in order for the limit to exist, the numerator must go to zero as well, otherwise the limit would be undefined

OpenStudy (anonymous):

ohhhh ok^

OpenStudy (anonymous):

therefore \[\lim_{h\to 0}f(a+h)-f(a)=0\] or \[\lim_{h\to 0}f(a+h)=f(a)\] therfore\(f\) is continuous at \(a\)

OpenStudy (anonymous):

as @jim_thompson5910 said, the converse if false

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

hmmmmm. @satellite73 But couldn't we argue the following? If a function is continuous at a, then:\[\lim_{h \rightarrow 0}f(a+h)=f(a)\]So,\[\lim_{h \rightarrow 0}f(a+h)-f(a)=0\]dividing both sides by h:\[\lim_{h \rightarrow 0} \frac{f(a+h)-f(a)}{h}\]if this limit exists. But this limit does exist since f(a+h) approaches f(a) as h goes to zero (it is continuous). So, the function is differentiable if it is continuous?

OpenStudy (anonymous):

you could argue that, but it would be wrong

OpenStudy (anonymous):

ok. why?

OpenStudy (anonymous):

because in general it is possible that you have \(\lim_{x\to a} g(x)=0\) and \(\lim_{x\to a}f(x)=0\) but \[\lim_{x\to a}\frac{f(x)}{g(x)}\] does not exist

OpenStudy (anonymous):

ah...ok. I would love to know why!

OpenStudy (anonymous):

getting \(\frac{0}{0}\) as a limiting form does not guarantee that the limit exists however getting \(\frac{b}{0}\) as a limiting form if \(b\neq 0\) guarantees that the limit does not exist

OpenStudy (anonymous):

hmmmm yeah that makes sense actually. Thinking of l'Hospital's rule for indeterminate forms and yeah I remember that some couldn't be resolved...

OpenStudy (anonymous):

the absolute value provides the canonical example if \(f(x)=|x|\) then \(\lim_{x\to 0}|x|=0\)

OpenStudy (anonymous):

but \(\lim_{x\to 0}\frac{|x|}{x}\) does not exist

OpenStudy (anonymous):

piecewise functions make a mess usually

OpenStudy (anonymous):

yeah...makes sense.

OpenStudy (anonymous):

^ agreed, I despise piecewise functions.

OpenStudy (anonymous):

I recall the abs value function and the kink in it at x=0. interesting stuff

OpenStudy (anonymous):

yeah they are a pain to write also of course in real life (if there is such a thing) almost all functions are piecewise

OpenStudy (anonymous):

lol